OpenStudy (anonymous):

if x+6=3sin(theta) for 0

OpenStudy (anonymous):

i replace theta by k for ease sin(k)=(x+6)/3 use the relation cos(2k)=cos^2(k)-sin^2(k)=1-2sin^2(k)

OpenStudy (anonymous):

it says that's incorrect no idea how If x + 6 = 3 \sin(\theta) for 0 < \theta < \pi/2 express \cos(2\theta) in terms of x. that's the full question